OpenStudy (anonymous):

‎37 , 41 , ? , 47 , 51 find?

OpenStudy (anonymous):

opts are. a) 43 b) 44 c) 45 d) 46

OpenStudy (anonymous):

how did u get that? its not in options.

OpenStudy (anonymous):

i got 43 but am not entirely sure.

OpenStudy (m):

doesnt make sense are you sure those are the numbers

OpenStudy (anonymous):

yeah....its correct question/

OpenStudy (anonymous):

at last its +4 not +2.

OpenStudy (m):

are they asking for pattern? is this statistics?

OpenStudy (anonymous):

they are from series!

OpenStudy (anonymous):

well my guess is like this....37+6=43, and 43+8=51 similarly,41+6=47,........seeing the the terms alternatively.

OpenStudy (m):

i guess that can work sorry not really familiar with series

OpenStudy (anonymous):

bingo i got an answer. its 43.

OpenStudy (anonymous):

cuz all of them are prime numbers.

OpenStudy (m):

ah i see gj!

OpenStudy (anonymous):

it should be 44. because 37+4=41; 41+3=44 44+47; 47+4=51;